OpenStudy (anthonyym):

Can I prove a proportionality equation by cross multiplying? If yes, then can I prove a trig identity that is a proportionality equation using cross multiplication? My teacher said we cannot "work across the equals sign"

OpenStudy (skullpatrol):

You are not "proving" a proportion by cross multiplying. Instead what you are doing is solving an equation to get a solution set. With trig identities you are proving that one side is the same as the other, so you can't cross multiply.

OpenStudy (anthonyym):

So if I have \[\frac{ 1 }{ 2 }=\frac{ 4 }{ 8 }\] I cannot prove this proportion by writing 1*8 = 4*2 8 = 8

OpenStudy (skullpatrol):

Yes, that is one way to show that they are equal. But you are dealing with only numerals and not variables, right?

OpenStudy (anthonyym):

Yes

OpenStudy (skullpatrol):

With variables you have to be careful that no denominator equals 0.

OpenStudy (anthonyym):

But you wouldn't write the proportion to prove if the denominator was already 0. I wouldn't write 1/0 = 4/8 and say prove the identity if it's already divided by 0.